OpenStudy (anonymous):

how do you find the inverse of f(x) = 2^x

OpenStudy (rulnick):

log_2 (x)

sam (.sam.):

y=2^x

sam (.sam.):

ln(y)=xln2 x=ln(y)/ln(2) \[f^{-1}(x)=\ln(x)/\ln(2)\]

OpenStudy (rulnick):

The inverse function for 2^x is log base 2 of x (log_2 (x)).

OpenStudy (rulnick):

... which can be written in terms of other logarithms, like ln, by using log_2 (x) = ln(x) / ln(2), as in what .Sam. wrote, for example. But the simplest answer is log_2 (x).